From PREX Wiki
WORK IN PROGRESS
- Please don't worry if there is a needless test on here, or an important one missing, or bad prioritization.
- Lets fix it
Optics and Q2
- Energy Measurement (Doug) (non-invasive)
- Priority High
- look for 1e-3 accuracy. Non invasive.
- Q2 measurement
- Repeat of Q2 measurement, with improved pedestal noise control.
- Detector counting mode.
- S0 trigger
- production raster and beam position
- Optics Data
- retake watercell,
- retake optimization data
- inelastic spectrum data
- Stubby Quartz / Blinded Quartz?
- Longish access to setup and back out, but short data taking period (production conditions).
- Replace detector quartz with superelastic stub or blinding cap (alternate US/DS and L/R)
- Integrating measurement. Could normalize increased gain just by lowering beam current in production condition.
- If surprising results found, continue test or plan second test
- Rescattering measurement (~1 shift)
- Detailed Rescattering Runplan
- perhaps plan for late april?
- repeat of previous measurements, but new optics and a crucial cross-check
- Detune spectrometer up in momentum, to dump elastic peak into spectrometer wall.
- Measure (integrating mode) signal in main detectors. Requires HV boot-strap to get higher gain (is 10x possible?)
- should scale all spectrometer magnets.
- At each dipole setting, also take data with nominal septum
- Once the elastic peak is out of the focal plane, it should be possible turn on the S3 and S0 scintillator. This should be done when possible, and the T1/T2/T3 trigger rate recorded for each run.
- Pole tip re-scattering
- Requires investigation or planning - do we have an idea here?
- Q2 or septum mis-tune, to increase interception on pole tip and improve poletip rescattering estimate
- A_T detectors
- Do we want to blind them again?
- Linearity Studies (PITA + current scans) (Caryn)
- These are ongoing - should continue regularly. ~ 1 hour every 3 days.
- Linearity LED tests
- This will be studied off line
- Sign check
- run Moller, Compton, and parity DAQ for quick PITA scan
- Have we already done this? Parasitic with any PITA scan
- Repeat measurements
- about ~30-60C. Goal: 1-1.5 shifts per measurement.
- Q2 scan?
- deadtime scan?
- target position scan?
- No-target runs
- Laser polarization optimization
- Laser polarization flip
- linearity studies